OpenStudy (anonymous):

Simplify and combine. (25x^2y^4)^3/2

OpenStudy (anonymous):

\[(25x^2y^4)^2/^3\] It's all raised to 2/3.

OpenStudy (anonymous):

\[(25x^2y^4)^{\frac{2}{3}}=25^{\frac{2}{3}}*x^{2*\frac{2}{3}}*y^{4*\frac{2}{3}}\] 3/2 would be much easier

OpenStudy (anonymous):

\[25^{\frac{3}{2}}=\sqrt{25}^3\]

OpenStudy (anonymous):

Wouldn't you multiply x^2 and y^4 by 3? so it would be essentially \[\sqrt{25^3x^6y^{12}}\]

OpenStudy (anonymous):

I think the answer is 125x^3y^6

OpenStudy (anonymous):

yes it is, and it would be easier, but some people cant actually understand that part

OpenStudy (anonymous):

Woo! I understood something! Thank you!